# Effects of dynamical friction on perturbations for evolving dark energy

### v1

@@ VI Effects of the modified evolution of the mass function
In this section we build upon results of previous sections and we investigate the halo mass function. This is an important observable, used to probe cosmology at late times by counting observed structures in the Universe. Differently from other observables, such as the CMB or the matter power spectrum, it is not model-independent, in the sense that an observed halo mass function relies on an assumed cosmology via the evaluation of the cosmic volume. This cosmology-dependence may be alleviated or eliminated considering observations at $z\simeq 0$.
@@
From a theoretical point of view, the functional form of the halo mass function can be inferred from fits to $N$-body simulations (Jenkins et al., 2001; Tinker et al., 2008; Watson et al., 2013), even though analytical derivations exist (Press and Schechter, 1974; Sheth et al., 2001). In this work we make a very conservative choice and use the Sheth-Tormen (ST) halo mass function (Sheth et al., 2001). This choice is dictated by the fact that Creminelli et al. (2010) used this formulation to describe the halo mass function in clustering dark energy models with null sound speed. We will come back to this point later.
The functional form of the halo mass function for Sheth et al. (2001) is
@@
where $\nu=\delta_{\rm c}/\sigma$, with $\sigma$ the variance of the linear matter power spectrum and $A=0.3222$, $a=0.75$ and $p=0.3$.
@@
We obtain the linear matter power spectrum from a suitably modified version of the Einstein-Boltzmann code class [^8] https://github.com/lesgourg/class_public/ Blas et al. (2011) [^9] For simplicity, we perform calculations in the Newtonian gauge.. For internal consistency, the code always evolves the perturbations of every species, therefore, to distinguish between smooth and clustering dark energy we set cs2_fld, the variable representing $c_{\rm s}^{2}$, to $1$ and 00, respectively.
Figure 6: Evolution of the matter power spectrum $P(k)$ at $z=0$ for the $\Lambda$CDM model. In the top (bottom) panel we show results for dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
@@
Unsurprisingly, dynamical friction is negligible on all the scales considered, except for the case with the largest $\eta_{0}$ and on very small scales. In this case, differences range between $5\%$ and $10\%$. Dissipative friction, instead, as already discussed before, shows a stronger effects, almost constant along the scales shown and with differences of about 25% with respect to the case without friction. As expected, differences point towards a lack of power, sign of a slowing down in the evolution of dark matter perturbations. This is due to the fact that we considered that all models have the same initial conditions, in particular the primordial amplitude of perturbations $A_{\rm s}$. This translates into a lower value of the normalization of the matter power spectrum today, $\sigma_{8}$.
@@
For dark energy models we find qualitative and quantitative similar results, therefore we will not show additional plots. Dynamical friction has a very small effect, while dissipative pressure friction is substantially more relevant. The exact quantitative effect depends on the particular dark energy model, but its behaviour is totally analogous to what seen for the growth factor, presented in Figs. 2 and 3. As perturbations in dark energy models are very small, even when $c_{\rm s}^{2}=0$, we do not see appreciable differences between smooth and clustering dark energy models.
@@
We can now discuss the effects of friction on the halo mass function. We present our results for the number of halos above a given mass for the $\Lambda$CDM model in Fig. 7 and for smooth dark energy models in Figs. 8 and 9.
@@
Figure 7: Number of objects above a given mass $M$ for the $\Lambda$CDM model at $z=0$. The top panel shows the effects of dynamical friction while the bottom panel refers to dissipative pressure friction. Line styles and colours are as in Fig. 1.
Dynamical friction has a small effect on the overall mass function and becomes appreciable only for the most extreme case with $\eta_{0}=0.01$. In this case, for masses of the order of $7\times 10^{14}\,M_{\odot}$, we notice a decrease of about 20% in the cumulative number of objects. On the same scales, the effect of dissipative pressure friction is about three times higher. As we evaluated the halo mass function today, this is a consequence of the lack of power highlighted in the matter power spectrum. We also remind the reader that for the $\Lambda$CDM model, the effect of $\Gamma$ is irrelevant. This result strengthens the previous discussion on the possibility of putting tight constraints on these models.
@@
Figure 8: Number of objects above a given mass threshold $M$ for the dark energy models studied in this work. We considered dynamical friction. Line styles and colours are as in Fig. 1.
Figure 9: As in Fig. 8, but for dissipative pressure friction.
@@
For the smooth case, all the models have the same qualitative behaviour when considering dynamical friction and the differences with respect to a frictionless model are largely independent of the particular equation of state used. As expected, differences increase at higher masses and for all the cases here considered we observe a decrease in the number of objects which form. This is coherent with the fact that friction dampens perturbations. The higher is $\eta_{0}$, the larger are the differences. For $\eta_{0}=10^{-4}$, deviations from the standard case where dynamical friction is absent, are, once again, negligible. Differences become appreciable for $\eta_{0}=10^{-3}$ and reach a value of about 2%–3% for masses of the order of $10^{15}\,M_{\odot}\,h^{-1}$, to reach 20%–25% for $\eta_{0}=10^{-2}$.
@@
More interesting is the case of dissipative pressure friction. Differences are enhanced with respect to the case with $\eta_{2}=0$ and can reach up to 75% of less objects. A particular case is model 6, whose equation of state is much smaller than $-1$. For this model, the suppression of structures is substantial, of the order of $50\%$ already for objects above $10^{13}\,M_{\odot}\,h^{-1}$.
@@
The analysis shows another interesting point. Whilst the behaviour of dynamical friction is coherent among the models, and percentage differences are roughly the same, for dissipative pressure friction this is not the case any more. First of all the exact difference strongly depends on the background equation of state. In fact, excluding model 6, we observe that differences range from 50% to 75% for masses of $10^{15}\,M_{\odot}\,h^{-1}$. But more importantly, analogously to what happens for the growth factor, we see that the two cases with $\Gamma=-1$ and $\Gamma=-2$ have a different behaviour with respect to each other: in fact, for models with a quintessence behaviour (models 1 and 3), the model with $\Gamma=-2$ predicts more objects than when $\Gamma=-1$. Opposite situation when the model is always or at least for some time in the phantom regime. Thus, the lower is the equation of state, the stronger are the differences.
VII Outlooks
@@
In this paper, we investigated the effects of friction on large scale structures. In this respect, we focused on both smooth and clustering dark energy models and limited our analysis to general relativity and to models with a linearly evolving equation of state, as these are more in line with observational data.
@@
We analysed the effects of friction on both linear and nonlinear perturbation regimes, employing the foundational principles of spherical collapse theory. To do so, we studied the evolution of perturbations and we determined the appropriate initial conditions to develop such perturbations.
Our findings revealed that the most significant discrepancies occur in the second hierarchical model, as it affects severely both linear and nonlinear perturbations. In this respect, we detailed the effects and consequences of friction on large-scale structures, highlighting the key differences from standard cases that do not incorporate such effects from the outset.
Further, we developed the consequences of the modified evolution of the cumulative mass function into the two aforementioned schemes. Consequences and discussions towards the underlying effects are reported, showing that our expectations are in line with current comprehension of our data. We demonstrated that the first recipe, dynamical friction proportional to the Hubble function, affects the mass function by 5% at most, while for the second model differences can be up to 75% for the range of masses probed in this work.
We also considered how perturbations of the dark energy affect the halo mass function. It turns out that the effects of clustering of dark energy is significantly more important than friction and the contribution of perturbations are strongly dependant on the background equation of state.
Consequently, with the severe exception of the second background treatment, our results seemed to suggest that there is no a priori reason to exclude the presence of friction, indicating that dissipation can potentially have a significant limiting effect on clustering.
For future works, we intend to extend our effort by involving more complicated friction models and also investigating more deeply the consequences on structure formation. Additional studies will be performed in showing how dark energy clustering can also be influenced by dark energy and/or we will clarify whether a possible interaction between the dark sector and friction may occur.
@@ Acknowledgements
F.P. acknowledges partial support from the INFN grant InDark, the Departments of Excellence grant L.232/2016 of the Italian Ministry of Education, University and Research (MUR) and from the European Union – Next Generation EU. FP also acknowledges the FCT project with ref. number PTDC/FIS-AST/0054/2021. F.P. also acknowledges the FCT project with ref. number PTDC/FIS-AST/0054/2021. The work of O.L. is partially financed by the Ministry of Education and Science of the Republic of Kazakhstan, Grant: IRN AP19680128.
@@ Appendix A Dark energy
In the next two subsections we will present results for the dark energy models considered in this work. We consider both smooth and clustering dark energy, but we will show figures about smooth dark energy models only, as the effect of clustering dark energy is very small and qualitatively we obtain the same results in both cases.
Figure 10: Evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ for the wCDM models studied here for the dynamical friction. Line styles and colours are as in Fig. 1.
Figure 11: As in Fig. 10, but for the dissipative pressure friction.
@@ A.1 Smooth dark energy
For smooth dark energy, we can infer analogous conclusions to the $\Lambda$CDM model. The effects of dynamical friction on the linearly extrapolated overdensity parameter $\delta_{\rm c}$ (Fig. 10) are small and of the same order of magnitude of the $\Lambda$CDM model and deviations from the case with $\eta_{0}=0$ are largely independent of the particular model considered.
For dissipative pressure friction (Fig. 11), as for the $\Lambda$CDM model, the overall amplitude of $\delta_{\rm c}$ is lower than when friction is not considered. The strength of the effect depends on the particular dark energy model. Models 2, 5 and 6 show the strongest deviations, while for the other three models differences are less pronounced. Differently from before, now the particular value of $\Gamma$ affects the final behaviour of the model. In fact, for the phantom model 6, when $\Gamma=-2$, the effect of friction is so large that the initial overdensity needs to reach unrealistic values so that perturbations can collapse. For this reason we do not show the results obtained. An indication of this comes from the bottom right panel of Fig. 3, where we see substantial differences with the standard case also at the level of the growth factor.
Regarding the effects of friction on the evolution of the virial overdensity $\Delta_{\rm vir}$, the results obtained are in line with what found for $\Lambda$CDM. In Fig. 12 we present our results for dynamical friction and in Fig. 13 for dissipative pressure friction. Also for dark energy models, dynamical friction has a very modest impact with differences of the same order of magnitude found for the $\Lambda$CDM model. Dissipative pressure friction has obviously a much stronger effect and we see an increase of the value of the virial overdensity. For models 2 and 5, in fact, the virial overdensity first increases and then decreases. It is also interesting to see that for model 3, the virial overdensity is almost constant over all the redshifts investigated.
Figure 12: Evolution of the virial overdensity $\Delta_{\rm vir}$ for the six dark energy models investigated with dynamical friction. Line styles and colours are as in Fig. 1.
Figure 13: As in Fig. 12, but for dissipative pressure friction.
Figure 14: Number of objects above a given mass threshold $M$ for the dark energy models studied in this work, when dark energy is perturbed. We considered dynamical friction. Line styles and colours are as in Fig. 1.
Figure 15: As in Fig. 14, but for dissipative pressure friction.
@@ A.2 Clustering dark energy
In this section we discuss results for clustering dark energy models, where we allow for perturbations of dark energy. Their strength is related to the value of the effective sound speed, defined, as previously discussed, by $c^{2}_{\rm eff}=\delta P_{\rm de}/\delta\rho_{\rm de}$. To fix it, one would need to consider a scalar field and its Lagrangian. We will, therefore, consider it as a free parameter. Quintessence models have $c^{2}_{\rm eff}=1$ and their perturbations are negligible on the scales of interest Unnikrishnan et al. (2008). For $k$-essence models we can achieve smaller values and in this work we will consider $c^{2}_{\rm eff}=0$ for which dark energy perturbations are larger and the dark energy fluctuations behave like matter.
Once again, we will consider the linear (represented by the growth factor) and the nonlinear evolution (represented by $\delta_{\rm c}$ e $\Delta_{\rm vir}$), but we will not present any figure, as there are no interesting nor special features even in this case.
In the linear regime, dark energy perturbations are always very subdominant with respect to matter perturbations and the evolution of $\delta_{\rm m}$ is totally unaffected by them. Similar results are obtained when considering the evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ and of the virial overdensity $\Delta_{\rm vir}$. We did not find appreciable differences with respect to the smooth case.
When dark energy is clustering, there is always the question of what is the correct definition of the density perturbation to be used. So far we have assumed that it is enough to consider dark energy affecting matter perturbations via modification of the Poisson equation and we found that dark energy perturbations do not affect matter perturbations appreciably. One could also assume that what really matters is not matter density perturbations, but rather the total density perturbation, defined as:
@@
Our results are presented in Fig. 16 for the case of dynamical friction and in Fig. 17 for dissipative pressure friction. We can now appreciate that the behaviour of $\delta_{\rm c,tot}$ is rather different from before and it is strongly dependant on the particular parameterisation of the equation of state used. Considering the effect of dynamical friction, we observe for models 2, 5 and 6 the typical behaviour obtained before. The linearly extrapolated overdensity approaches the EdS value (dark energy perturbations are negligible and furthermore scaled by $\Omega_{\rm de}$) and decreases at late times. For the other models, instead, we recover the EdS value at early time but at late times $\delta_{\rm c,tot}$ grows, showing that dark energy perturbations can have an appreciable effect. It is also interesting to notice that the effective value of dynamical friction becomes rather irrelevant, as all the models behave similarly to the case of no friction with $\eta_{0}=0$.
More complex is the analysis for dissipative pressure friction. Results do not only depend on the equation of state used, but also on the particular value of the parameter $\Gamma$. We thus observe a similar trend to what found in Fig. (16). Models 2, 5 and 6 behave as one would expect for smooth dark energy or for associating $\delta_{\rm c}$ to matter only. The other three models, instead, have an EdS-like behaviour at early times and grow, rather than decrease, at late times. It is also interesting to see that the dependence on the exponent $\Gamma$ becomes very shallow.
Figure 16: Evolution of the total linearly extrapolated overdensity $\delta_{\rm c}$ for the wCDM models studied here for dynamical friction. Line styles and colours are as in Fig. 1.
Figure 17: Evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ for the wCDM models studied here for dissipative pressure friction. Line styles and colours are as in Fig. 1.
This discussion highlights an important point about the correct definition of the density perturbations, as, even though $\delta_{\rm de}$ is smaller than $\delta_{\rm m}$ in general and have limited effects on the evolution of the latter, it can substantially affect our observables if we use $\delta_{\rm tot}$ rather than $\delta_{\rm m}$, as it was also demonstrated, discussing virialization, in Pace and Schimd (2022).

### v2

@@ 6 Effects of the modified evolution of the mass function
In this section, we build upon results of previous sections, and we investigate the halo mass function. This is an important observable that is used to probe cosmology at late times by counting isolated bounded halos in the Universe. As in the case of the CMB or the matter power spectrum, the halo mass function depends on the cosmological model assumed.
@@
From a theoretical point of view, the functional form of the halo mass function can be inferred from fits to $N$-body simulations [95, 96, 97], although analytical derivations exist [98, 99]. In this work, we make a very conservative choice and use the Sheth-Tormen (ST) halo mass function [99]. This choice is dictated by the fact that [100] used this formulation to describe the halo mass function in clustering dark energy models with null sound speed. We will come back to this point later.
The halo mass function is
$\frac{\mathrm{d}n}{\mathrm{d}M}=\frac{\bar{\rho}_{\rm,0}}{M^{2}}f(\nu)\left|\frac{\mathrm{d}\ln{\sigma}}{\mathrm{d}\ln{M}}\right|\,,$ (38)
where $f(\nu)$ is the multiplicity function, for which we use the following functional form [99]
@@
where $\nu=\delta_{\rm c}/\sigma$, with $\sigma$ the variance of the linear matter power spectrum and $A=0.3222$, $a=0.75$ and $p=0.3$ as best fit parameters of the theoretical multiplicity function compared to the one obtained numerically [99].
@@
We obtain the linear matter power spectrum from a suitably modified version of the Einstein-Boltzmann code class [^10] https://github.com/lesgourg/class_public/ [101]. For internal consistency, the code always evolves the perturbations of every species, therefore, to distinguish between smooth and clustering dark energy, we set cs2_fld, the variable representing $c_{\rm s}^{2}$, to $1$ and 00, respectively [^11] For simplicity, we perform calculations in the Newtonian gauge..
Figure 5: Evolution of the matter power spectrum $P(k)$ at $z=0$ for the $\Lambda$CDM model. In the left (right) panel we show results for dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
@@
Dynamical friction is negligible on all the scales considered, except for the case with the largest $\eta_{0}$ and on very small scales. In this case, the differences range between $5\%$ and $10\%$. Dissipative friction, instead, as already discussed, shows stronger effects, almost constant along the scales shown and with differences of about 25% with respect to the case without friction. As expected, differences point towards a lack of power, sign of a slowing down in the evolution of dark matter perturbations. This is due to the fact that we considered that all models have the same initial conditions, in particular the primordial amplitude of perturbations $A_{\rm s}$. This translates into a lower value of the normalization of the matter power spectrum today, $\sigma_{8}$.
@@
For dark energy models, we find qualitative and quantitative similar results; therefore, we will not show additional plots. Dynamical friction has a very small effect, whereas dissipative pressure friction is substantially more relevant. The exact quantitative effect depends on the particular dark energy model, but its behavior is totally analogous to what is seen for the growth factor, presented in Fig. 2. As perturbations in dark energy models are very small, even when $c_{\rm s}^{2}=0$, we do not see appreciable differences between smooth and clustering dark energy models.
@@
We can now discuss the effects of friction on the halo mass function. We present our results for the number of halos above a given mass for the $\Lambda$CDM model in Fig. 6 and for the smooth dark energy model DE1 in Figs. 7.
@@
Figure 6: Number of objects above a given mass $M$ for the $\Lambda$CDM model at $z=0$. The left panel shows the effects of dynamical friction while the right panel refers to dissipative pressure friction. Line styles and colours are as in Fig. 1.
@@
Figure 7: Number of objects above a given mass threshold $M$ for the dark energy model DE1 with respect to the $\Lambda$CDM prediction. We considered dynamical friction (left panel) and dissipative pressure friction (right panel). Line styles and colours are as in Fig. 1.
Dynamical friction has a small effect on the overall mass function and becomes appreciable only for the most extreme case with $\eta_{0}=10^{-2}$. In this case, for masses of the order of $7\times 10^{14}\,M_{\odot}\,h^{-1}$, we notice a decrease of about 20% in the number of objects with $M>10^{15}\,h^{-1}\,M_{\odot}$ for $\eta_{0}=0.01$, but also a decrease of roughly 10% for objects with masses similar to galaxy groups. For lower values of $\eta_{0}$, differences with respect to $\Lambda$CDM are much smaller and probably difficult to measure. These values are approximately the same for all six models. On the same scales, the effect of dissipative pressure friction is about three times higher. As we evaluated the halo mass function today, this is a consequence of the lack of power highlighted in the matter power spectrum. We also remind the reader that for the $\Lambda$CDM model, the effect of $\Gamma$ is irrelevant. This result strengthens the previous discussion on the possibility of placing tight constraints on these models.
@@
For the smooth case, all the models have the same qualitative behavior when considering dynamical friction and the differences with respect to a frictionless model are largely independent of the particular equation of state used. As expected, differences increase at higher masses and for all the cases here considered we observe a decrease in the number of objects which form. This is consistent with the fact that friction dampens perturbations. The higher $\eta_{0}$, the larger the differences. For $\eta_{0}=10^{-4}$, deviations from the standard case where dynamical friction is absent are again negligible. The differences become appreciable for $\eta_{0}=10^{-3}$ and reach a value of about 2%–3% for masses of the order of $10^{15}\,M_{\odot}\,h^{-1}$, to reach 20%–25% for $\eta_{0}=10^{-2}$.
@@
More interesting is the case of dissipative pressure friction. The differences are enhanced with respect to the case with $\eta_{2}=0$ and can reach up to 75% fewer objects. A particular case is model DE6, whose equation of state is much smaller than $-1$. For this model, the suppression of structures is substantial, of the order of $50\%$ already for objects above $10^{13}\,M_{\odot}\,h^{-1}$.
@@
The analysis shows another interesting point. Whilst the behavior of dynamical friction is coherent among the models, and percentage differences are roughly the same, for dissipative pressure friction this is not the case any more. First of all, the exact difference strongly depends on the background equation of state. In fact, excluding model DE6, we observe that differences range from 50% to 75% for masses of $10^{15}\,M_{\odot}\,h^{-1}$. But more importantly, analogously to what happens for the growth factor, we see that the two cases with $\Gamma=-1$ and $\Gamma=-2$ have a different behavior with respect to each other: in fact, for models with quintessence behavior (models DE1 and DE3), the model with $\Gamma=-2$ predicts more objects than when $\Gamma=-1$. Opposite situation where the model is always, or at least for some time, in the phantom regime. Thus, the lower the equation of state, the stronger the differences.
7 Outlooks
@@
In this paper, we investigate the effects of friction on large-scale structures. In this respect, we focussed on both smooth and clustering dark energy models and limited our analysis to general relativity and to models with a linearly evolving equation of state, as these are more in line with observational data.
@@
We analyze the effects of friction on both linear and nonlinear perturbation regimes employing the foundational principles of spherical collapse theory. To do so, we studied the evolution of perturbations, and we determined the appropriate initial conditions to develop such perturbations.
Our findings revealed that the most significant discrepancies occur in the second hierarchical model, as it severely affects both linear and nonlinear perturbations. In this respect, we detailed the effects and consequences of friction on large-scale structures, highlighting the key differences from standard cases that do not incorporate such effects from the outset.
In addition, we developed the consequences of the modified evolution of the cumulative mass function in the two aforementioned schemes. Consequences and discussions towards the underlying effects are reported, showing that our expectations are in line with current comprehension of our data. We demonstrated that the first recipe, dynamical friction proportional to the Hubble function, affects the mass function by 5% at most, while for the second model differences can be up to 75% for the range of masses invesigated in this work.
We also considered how perturbations of dark energy affect the halo mass function. It turns out that the effects of clustering of dark energy are significantly more important than friction and that the contributions of perturbations are strongly dependent on the background equation of state.
Consequently, with the severe exception of the second background treatment, our results seem to suggest that there is no a priori reason to exclude the presence of friction, indicating that dissipation can potentially have a significant limiting effect on clustering.
In future work, we intend to extend our effort by involving more complicated friction models and also investigating the consequences on structure formation. Additional studies will be performed to show how dark energy clustering can also be influenced by dark energy, and/or we will clarify whether a possible interaction between the dark sector and friction may occur.
CRediT authorship contribution statement
Francesco Pace: Supervision, Project administration. Orlando Luongo: Writing – review & editing, Writing – original draft, Investigation. Antonino Del Popolo: Validation, Conceptualization.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgements
F.P. acknowledges partial support from the INFN grant InDark and from the Italian Ministry of University and Research (mur), PRIN 2022 ‘EXSKALIBUR – Euclid-Cross-SKA: Likelihood Inference Building for Universe’s Research’, Grant No. 20222BBYB9, CUP C53D2300131 0006, and from the European Union – Next Generation EU. FP also acknowledges support from the FCT project “BEYLA – BEYond LAmbda” with ref. number PTDC/FIS-AST/0054/2021.
@@ Appendix A Dark energy
In the next two subsections we will present results for the dark energy models considered in this work. We consider both smooth and clustering dark energy, but we will show figures about smooth dark energy models only, as the effect of clustering dark energy is very small, and qualitatively we obtain the same results in both cases.
Figure 8: Evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ for the model DE1 for dynamical friction (left panel) and dissipative pressure friction (right panel). Line styles and colours are as in Fig. 1.
@@ A.1 Smooth dark energy
For smooth dark energy, we can infer conclusions analogous to the $\Lambda$CDM model. The effects of dynamical friction on the linearly extrapolated overdensity parameter $\delta_{\rm c}$ (left panel of Fig. 8) are small and of the same order of magnitude as the $\Lambda$CDM model and deviations from the case with $\eta_{0}=0$ are largely independent of the particular model considered.
For dissipative pressure friction (right panel of Fig. 8), as for the $\Lambda$CDM model, the overall amplitude of $\delta_{\rm c}$ is lower than when friction is not considered. The strength of the effect depends on the particular dark energy model. Models DE2, DE5 and DE6 show the strongest deviations, while for the other three models the differences are less pronounced. Unlike before, now the particular value of $\Gamma$ affects the final behavior of the model. In fact, for the phantom model DE6, when $\Gamma=-2$, the effect of friction is so large that the initial overdensity must reach unrealistic values so that the perturbations can collapse. For this reason, we do not show the results obtained. This is the same behavior we observed when studying the growth factor.
Regarding the effects of friction on the evolution of virial overdensity $\Delta_{\rm vir}$, the results obtained are in line with what found for $\Lambda$CDM. In the left (right) panel of Fig. 9 we present our results for dynamical friction (dissipative pressure friction). Also for dark energy models, dynamical friction has a very modest impact with differences of the same order of magnitude found for the $\Lambda$CDM model. Dissipative pressure friction obviously has a much stronger effect, and we see an increase in the value of the virial overdensity. For models DE2 and DE5, in fact, the virial overdensity first increases and then decreases. It is also interesting to see that for model DE3, the virial overdensity is almost constant over all the redshifts investigated.
Figure 9: Evolution of the virial overdensity $\Delta_{\rm vir}$ for the model DE1 for dynamical friction (left panel) and dissipative pressure friction (right panel). Line styles and colours are as in Fig. 1.
@@ A.2 Clustering dark energy
In this section, we discuss results for clustering dark energy models, where we allow for perturbations of dark energy. Their strength is related to the value of the effective sound speed, defined, as previously discussed, by $c^{2}_{\rm eff}=\delta P_{\rm de}/\delta\rho_{\rm de}$. To fix it, one would need to consider a scalar field and its Lagrangian. We will therefore consider it as a free parameter. Quintessence models have $c^{2}_{\rm eff}=1$ and their perturbations are negligible on the scales of interest [102]. For $k$-essence models we can achieve smaller values and in this work we will consider $c^{2}_{\rm eff}=0$ for which dark energy perturbations are larger and the dark energy fluctuations behave like matter.
Once again, we will consider the linear (represented by the growth factor) and the nonlinear evolution (represented by $\delta_{\rm c}$ and $\Delta_{\rm vir}$), but we will not present any figure, as there are no interesting nor special features even in this case.
In the linear regime, dark energy perturbations are always very subdominant with respect to matter perturbations, and the evolution of $\delta_{\rm m}$ is totally unaffected by them. Similar results are obtained when considering the evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ and of the virial overdensity $\Delta_{\rm vir}$. We did not find appreciable differences with respect to the smooth case.
When dark energy is clustering, there is always the question of what is the correct definition of the density perturbation to be used. So far we have assumed that it is enough to consider dark energy affecting matter perturbations via modifications of the Poisson equation, and we found that dark energy perturbations do not affect matter perturbations appreciably. One could also assume that what really matters is not matter density perturbations, but rather the total density perturbation, defined as:
@@
Our results are presented in Fig. 10. We can appreciate that the behavior of $\delta_{\rm c,tot}$ is rather different from before and strongly depends on the particular parameterization of the equation of state used. Considering the effect of dynamical friction, we observe for models DE2, DE5 and DE6 the typical behavior obtained before. The linearly extrapolated overdensity approaches the EdS value (dark energy perturbations are negligible and furthermore scaled by $\Omega_{\rm de}$) and decreases at late times. For the other models, instead, we recover the EdS value at an early time but at late times $\delta_{\rm c,tot}$ grows, showing that dark energy perturbations can have an appreciable effect. It is also interesting to note that the effective value of dynamical friction becomes rather irrelevant, as all the models behave similarly to the case of no friction with $\eta_{0}=0$.
More complex is the analysis for dissipative pressure friction. The results depend not only on the equation of state used, but also on the particular value of the parameter $\Gamma$. Thus, we observe a similar trend to that found in Fig. 10. Models DE2, DE5 and DE6 behave as one would expect for smooth dark energy or for associating $\delta_{\rm c}$ to matter only. The other three models, instead, have EdS-like behavior at early times and grow, rather than decrease, at late times. It is also interesting to see that the dependence on the exponent $\Gamma$ becomes very shallow.
Figure 10: Evolution of the total linearly extrapolated overdensity $\delta_{\rm c}$ for the model DE1 for dynamical friction (left panel) and dissipative pressure friction (right panel). Line styles and colours are as in Fig. 1.
This discussion highlights an important point about the correct definition of density perturbations, as although $\delta_{\rm de}$ is smaller than $\delta_{\rm m}$ in general and has limited effects on the evolution of the latter, it can substantially affect our observables if we use $\delta_{\rm tot}$ instead of $\delta_{\rm m}$, as was also demonstrated in discussing virialization in [53].